OpenStudy (anonymous):

The magnitude of the resultant of two equal forces is 4root3 kg.wt if the measure of the angle between the resultant is 30. find the two forces

OpenStudy (anonymous):

R² = A² + B² + 2ABcosθ

OpenStudy (anonymous):

R-resultant force A,B-the two forces theta-angle between them...

OpenStudy (anonymous):

 Ftot = 2*F*cos(30°) <-> F = Ftot / (2*cos(30°))

OpenStudy (anonymous):

are the two forces equal to 4 units and the angle between the tow forces are 60...